# Multi-Review Fusion-in-Context

### v1

@@ Abstract
Grounded text generation, encompassing tasks such as long-form question-answering and summarization, necessitates both content selection and content consolidation. Current end-to-end methods are difficult to control and interpret due to their opaqueness. Accordingly, recent works have proposed a modular approach, with separate components for each step. Specifically, we focus on the second subtask, of generating coherent text given pre-selected content in a multi-document setting. Concretely, we formalize Fusion-in-Context (FiC) as a standalone task, whose input consists of source texts with highlighted spans of targeted content. A model then needs to generate a coherent passage that includes all and only the target information. Our work includes the development of a curated dataset of 1000 instances in the reviews domain, alongside a novel evaluation framework for assessing the faithfulness and coverage of highlights, which strongly correlate to human judgment. Several baseline models exhibit promising outcomes and provide insightful analyses. This study lays the groundwork for further exploration of modular text generation in the multi-document setting, offering potential improvements in the quality and reliability of generated content.  [^1] Our benchmark, FuseReviews, including the dataset, evaluation framework and designated leaderboard, can be found at https://fusereviews.github.io/.
@@ 1 Introduction
To promote research on FiC, we start by formally defining the task (§3). We then introduce a dataset (§4), carefully constructed via controlled crowdsourcing (Roit et al., 2020). Each of its 1000 instances comprises a set of inputs with highlights, and a corresponding fused text. The dataset is created through an efficient procedure, adapted from (Slobodkin et al., 2022), leveraging existing multi-document summarization datasets, specifically in the business reviews domain. We also develop an evaluation framework (§5) that assesses outputs’ faithfulness and coverage of highlights. The dataset and evalation framework are released as a benchmark, called FuseReviews. We explore various baseline models on the benchmark and report their performance (§6). Our findings reveal that while these models show promising results, there is still room for further improvement in future research.
@@ 3 Task Definition
Next, we describe a dataset (§4) and an evaluation framework (§5) that comply with the task definition. These are released as the FuseReviews benchmark for the FiC task.
@@ 4 Dataset for FiC
To comply with the task definition, an instance in a FiC dataset is expected to be a document set $D$ with marked spans $H=\{h_{1},h_{2},...,h_{n}\}$, and a corresponding fused text $f$. To compile such data, we leverage existing multi-document summarization datasets and extract high-quality FiC instances via controlled crowdsourcing (Roit et al., 2020), following Slobodkin et al. (2022), while adapting their method to the multi-text setting, and the business reviews domain.
@@ 4.1 Dataset Collection
Given a document set $D$ and corresponding reference summary $\widehat{f}$ from an existing multi-document summarization dataset, the annotation process aims to identify the spans in the source texts $\{h_{1},h_{2},...,h_{n}\}$ that cover all the information in $\widehat{f}$. This approach simplifies the annotation process compared to annotating from scratch, i.e., reading documents, marking highlights according to some specifications, and writing a coherently-fused text, which is reminiscent of standard formation of multi-document summarization datasets. Conversely, our approach requires locating and aligning spans between the source text and the already available reference summary, essentially ‘‘backwards engineering’’ the original human summarization process.
@@
To facilitate the annotation of alignments between reviews and their corresponding summary, we adapt a web-based annotation tool from (Slobodkin et al., 2022), and deploy it on Amazon Mechanical Turk222www.mturk.com for crowdsourcing (§4.3 will explain the controlled crowdsourcing procedure). The application presents reviews and the respective summary side-by-side, and annotators are guided to highlight pairs of spans in the reviews and the summary that directly align. To reduce cognitive load, a summary is displayed alongside an individual review, and focus is placed on one summary sentence at a time. To further ease the process for annotators, lemmas in the review overlapping with lemmas in the currently focused summary sentence are emboldened. [^3] Lemmatizing with spaCy Honnibal and Montani (2017). This enables quick skimming through the review, however workers are trained not to rely solely on exact matches for highlighting (as reflected in §4.2 and §4.3).
@@
Annotators are instructed to align statements dealing with a single aspect of a hotel or business (e.g., ‘‘room cleanliness’’) from the summary with the most relevant spans in the reviews, and to do so for all aspects in the summary in order to cover the whole summary text. [^4] We observed that instructing annotators to focus on one aspect at a time enhances the efficiency in locating the relevant review spans, particularly when a summary sentence includes content that is scattered across different parts of the review. See §4.2 for the detailed annotation guidelines. This, in turn, creates instances of highlighted spans within reviews, with a corresponding coherent fusion of those highlights (the summary). Each review-summary pair is annotated by a single trained annotator. To enhance quality, submissions are randomly sampled and reviewed, with feedback provided as necessary.
@@
In total we sampled 1000 instances of review-set/summary pairs (700 instances from CocoTrip and 300 from FewSum). See Table 1 for full statistics. [^5] See Appendix F for more details.
@@ 4.2 Annotation Guidelines and Data Traits
Figure 2 illustrates the annotation flow. When presented with a review of an entity (hotel or business) and a summary with a sentence in focus, an annotator first identifies aspects of the entity within the focused summary sentence. An aspect is not simply a facet of an entity, such as ‘‘rooms’’ or ‘‘staff’’, but more specifically it is a characteristic, such as ‘‘room cleanliness’’, ‘‘room style’’ or ‘‘staff helpfulness’’. (More on this in Appendix A.1.)
@@
Additionally, reviews often express varying opinions about the same aspect, such as “the service was great” as opposed to “the staff was unprofessional”. When summarizing, all these varying opinions should be considered to reflect the overall sentiment. As a result, summary segments may range from statements like “the staff was overall liked” to “some people liked the staff”, depending on the spectrum of opinions. Hence, to properly capture this consolidation of differing viewpoints, annotators are also guided to align review mentions that either sentimentally entail or contradict the summary aspect. For example, the two above conflicting review spans should be aligned to the summary span “the service was mostly good”.
Finally, annotators may mark multiple spans in reviews that redundantly represent the same statement. The guidelines also address paraphrasing, non-consecutive highlights, and un-alignable summary spans. A detailed explanation of these guidelines is in Appendix A.2.
@@ 4.3 Annotator Training
The requirements of the aforementioned annotation process call for proficient-level annotations, which we achieved by means of controlled crowdsourcing Roit et al. (2020). We identified qualified annotators through three open qualification rounds, followed by three closed rounds for selected annotators, focusing on further training and refinement. Each open round involves annotators reading a brief task description and accordingly aligning information between a single summary sentence and a short review, on a simplified interface. After each open round we reviewed the alignment and provided feedback. We then checked whether the annotator implemented our feedback in the following round (with a different sentence-review instance). If the annotator satsfyingly cooperated throughout the open rounds, they moved on to the closed rounds. The annotator started by watching a 25-minute tutorial on the full annotation tool and guidelines (§4.1 and §4.2). The closed rounds were conducted similarly to the open rounds, but with a whole summary and review, with all guidelines, and on the full interface. The qualification process was fully compensated with a customary wage, requiring up to 5 minutes per round. From this process, we were able to gather 8 trained annotators, who annotated the 1000 instances in our dataset.
@@ 5.3 Trained Coverage Metric
Inspired by recent work that evaluates faithfulness and factuality using a dedicated trained model (Yin et al., 2021; Utama et al., 2022; Gekhman et al., 2023; Soleimani et al., 2023), we finetune an LLM that is tasked to assess whether the generated passage fully covers the highlights. In our methodology, each highlight is individually input along with the entire output, and the model outputs a binary answer for whether the highlight is contained in the passage. [^7] We also tried concatenating all the highlights together, and found it to be inferior.. We derive synthesized training data for this task from our FiC dataset, using highlights and their corresponding summaries. For negative samples, we remove the summary sentence that aligns with the highlight. For positive samples, a random non-aligning summary sentence is omitted (to avoid a potential bias caused by sentence exclusion in the negative samples). We finetune a flan-t5-large model (Chung et al., 2022) with the synthesized coverage data. The input to the model is the highlight and modified summary, and the output is ‘yes’ or ‘no’, for positive and negative samples, respectively. The final score is the average probability of the token ‘yes’ across all highlights. [^8] We also explored an NLI-based coverage metric, where the passage serves as the premise and the highlights function as the hypothesis. We found it to achieve comparable results, however it requires substantially more computation time and memory. For more details, see Appendix E.
@@ 6.1 Experimental Setup
We examine several baseline models for solving the FiC task. The input to a model is a document set with spans marked within the documents (highlights), and the model ouputs a fused passage around the highlights.
@@ 6.2 Results
We apply our evaluation metrics on the proposed systems, with results presented in Table 3. We first observe that the exclusion of context from the input (Flan-T5only-H) yields the strongest faithfulness and coverage scores, yet the lowest coherence and redundancy scores. This shows the importance of incorporating context for more seamless outputs. Meanwhile, the removal of highlights (Flan-T5no-H) leads to a substantial degradation in faithfulness and coverage. This indicates that the model indeed succeeds in learning to adjust the output according to the highlights, underlining the highlights’ role in enhancing the model’s performance.
Interestingly, even though the RL reward functions used in the RL-enriched model are the faithfulness and coverage metrics themselves, the outputs are eventually negatively affected when evaluating with these metrics. This result calls for a more in-depth investigation of enhanced reward functions that can leverage the benefits of RL-enrichment, as was shown to be helpful in Slobodkin et al. (2023a) for the single-input setup. We also find that single-shot GPT-4 yields the most coherent and least redundant texts. While it ranks highly in faithfulness and coverage, it is still overtaken by the finetuned Flan-T5only-H. Overall, our findings invite for further research on the FiC task, to develop fusion strategies that ensure comprehensive coverage and faithfulness to highlighted content, with coherent and low-redundancy outputs.
@@ 7 Conclusion
In this paper, we further promote the decomposition of grounded text generation as presented in (Slobodkin et al., 2022), extending it to the multi-document setting. To that end, we introduce the Fusion-in-Context (FiC) task, an extension of the task from (Slobodkin et al., 2022) which focuses on the content fusion step, to the multi-document setting. The FiC setting facilitates employing a single general-purpose fusion model for diverse content selection needs, capturing the challenges of repetitiveness and contradictions in source documents. It also supports interactive, user-driven generation, by allowing users to choose personalized content for the FiC module to merge into a customized passage. Moreover, the direct access to the pre-selected ‘‘highlights’’ facilitates attributed generation, where the pre-selected segments also serve as supporting cited content for the fused text. To advance the task, we prepare a
@@ Appendix B NLI Zero-Shot Prompt
Table 4: Faithfulness, coverage and F-1 scores of the zero-shot GPT-4 model on 30 instances for the FiC development set, for varying numbers of in-context examples in the prompt. For each metric, the best scores are in bold.
@@ Appendix C GPT-4 Prompting
Table 4 presents the faithfulness, coverage and F-1 scores of the zero-shot GPT-4 model across 30 instances from the FiC development set, for varying numbers of in-context examples in the prompt. Based on these outcomes, we chose to proceed with a single in-context example.
@@ Appendix D Fluency and Redundancy Human Annotation Protocol
We ask crowd-workers to assess the fluency and redundnacy of the texts produced by all models under examination. We employ annotators who have demonstrated proficiency in semantic tasks, including summarization, in previous experiments. For evaluation purposes, 100 instances are randomly selected from our test set, and the texts generated by each model for these instances are evaluated, resulting in 500 total samples. Each sample is reviewed by three different annotators, and their scores are averaged to obtain a final assessment. The evaluation is facilitated through two Amazon Mechanical Turk interfaces, specifically designed for this study. One interface focuses on evaluating coherence, while the other assesses redundancy, with each interface presenting the annotators with one of the 500 samples (as depicted in Figure 4). Consistent with the methodology of (Slobodkin et al., 2022), a 5-point Likert scale is employed to rate the fluency and redundancy of the generated summaries. To minimize ambiguity and promote consistent ratings, each score on the scale is accompanied by explicit criteria (also illustrated in Figure 4). Taking into account an average response time of 30 seconds for each evaluation, we set the compensation for each response at 10 ¢.
@@
Figure 4: Example of the data collection interfaces used by the crowd-workers to evaluate the fluency (3(a)) and redundancy (3(b)) of summaries.
@@ E.3 Additional Meta Evaluation Setup Details
To achieve further agreement between the three authors, a supplementary reconciliation procedure was undertaken for the ten instances annotated by all three authors. This procedure entailed discussions for each instance where the annotations diverged by more than one point, separately for the faithfulness and coverage scores. During these discussions, each author explained the rationale behind their assigned score. Subsequently, the authors endeavored to reach a unanimous agreement on each instance, thereby further aligned their scoring criteria.

### v2

@@ Abstract
Grounded text generation, encompassing tasks such as long-form question-answering and summarization, necessitates both content selection and content consolidation. Current end-to-end methods are difficult to control and interpret due to their opaqueness. Accordingly, recent works have proposed a modular approach, with separate components for each step. Specifically, we focus on the second subtask, of generating coherent text given pre-selected content in a multi-document setting. Concretely, we formalize Fusion-in-Context (FiC) as a standalone task, whose input consists of source texts with highlighted spans of targeted content. A model then needs to generate a coherent passage that includes all and only the target information. Our work includes the development of a curated dataset of 1000 instances in the reviews domain, alongside a novel evaluation framework for assessing the faithfulness and coverage of highlights, which strongly correlate to human judgment. Several baseline models exhibit promising outcomes and provide insightful analyses. This study lays the groundwork for further exploration of modular text generation in the multi-document setting, offering potential improvements in the quality and reliability of generated content. [^1] Our benchmark, FuseReviews[Uncaptioned image], including the dataset, evaluation framework, and designated leaderboard, can be found at https://fusereviews.github.io/.
@@ 1 Introduction
To promote research on FiC, we start by formally defining the task (§3). We then introduce a dataset (§4), carefully constructed via controlled crowdsourcing (Roit et al., 2020). Each of its 1000 instances comprises a set of inputs with highlights, and a corresponding fused text. The dataset is created through an efficient procedure, adapted from Slobodkin et al. (2022), leveraging existing multi-document summarization datasets, specifically in the business reviews domain. We also develop an evaluation framework (§5) that assesses outputs’ faithfulness and coverage of highlights. The dataset and evaluation framework are released as a benchmark, called FuseReviews[Uncaptioned image]. We explore various baseline models on the benchmark and report their performance (§6). Our findings reveal that while these models show promising results, there is still room for further improvement in future research.
@@ 3 Task Definition
Next, we describe a dataset (§4) and an evaluation framework (§5) that comply with the task definition. These are released as the FuseReviews[Uncaptioned image] benchmark for the FiC task.
@@ 4 Dataset for FiC
To comply with the task definition, an instance in a FiC dataset is expected to be a document set $D$ with marked spans $H=\{h_{1},h_{2},...,h_{n}\}$, and a corresponding fused text $f$. To compile such data, we leverage existing multi-document summarization datasets and extract high-quality FiC instances via controlled crowdsourcing (Roit et al., 2020), by adapting the method from Slobodkin et al. (2022) to the multi-text setting, and the business reviews domain.
@@ 4.1 Dataset Collection
Given a document set $D$ and corresponding reference summary $\widehat{f}$ from an existing multi-document summarization dataset, the annotation process aims to identify the spans in the source texts $\{h_{1},h_{2},...,h_{n}\}$ that cover all the information within $\widehat{f}$. This approach simplifies the annotation process compared to annotating from scratch, i.e., reading documents, marking highlights according to some specifications, and writing a coherently-fused text, which is reminiscent of standard formation of multi-document summarization datasets. Conversely, our approach requires locating and aligning spans between the source text and the already available reference summary, essentially ‘‘reverse engineering’’ the original human summarization process.
@@
To facilitate the annotation of alignments between reviews and their corresponding summary, we adapt a web-based annotation tool from Slobodkin et al. (2022), and deploy it on Amazon Mechanical Turk222www.mturk.com for crowdsourcing (§4.3 will explain the controlled crowdsourcing procedure). The application presents reviews and the respective summary side-by-side, and annotators are guided to highlight pairs of spans in the reviews and the summary that directly align. To reduce cognitive load, a summary is displayed alongside an individual review, and focus is placed on one summary sentence at a time. To further ease the process for annotators, lemmas in the review overlapping with lemmas in the currently focused summary sentence are emboldened. [^3] Lemmatizing with spaCy Honnibal and Montani (2017). This enables quick skimming through the review, however, workers are trained not to rely solely on exact matches for highlighting (as discussed in §4.2 and §4.3).
@@
Annotators are guided to align statements dealing with a single aspect of a hotel or business (e.g., ‘‘room cleanliness’’) from the summary with the most relevant spans in the reviews, and to do so for all summary aspects in order to cover the whole summary text (see §4.2 for the detailed annotation guidelines). [^4] We observed that instructing annotators to focus on one aspect at a time enhances the efficiency in locating the relevant review spans, particularly when a summary sentence includes content that is scattered across different parts of the review. This, in turn, creates instances of highlighted spans within reviews, with a corresponding coherent fusion of those highlights (the summary). Each review-summary pair is annotated by a single trained annotator. To enhance quality, submissions are randomly sampled and reviewed, with feedback provided as necessary.
@@
In total, we sampled 1000 instances of review-set/summary pairs (700 instances from CocoTrip and 300 from FewSum). See Table 1 for full statistics. [^5] See Appendix F for more details.
@@ 4.2 Annotation Guidelines and Data Traits
Figure 2 illustrates the annotation flow. When presented with a review of an entity (hotel or business) and a summary with a sentence in focus, an annotator first identifies aspects of the entity within the focused summary sentence. An aspect is not simply a facet of an entity, such as ‘‘rooms’’ or ‘‘staff’’, but more specifically it is a characteristic, such as ‘‘room cleanliness’’, ‘‘room style’’ or ‘‘staff helpfulness’’ (more on this in Appendix A.1).
@@
Additionally, reviews often express varying opinions about the same aspect, such as “the service was great” as opposed to “the staff was unprofessional”. When summarizing, all these varying opinions should be considered to reflect the overall sentiment. As a result, summary segments may range from statements like “the staff was overall liked” to “some people liked the staff”, depending on the spectrum of opinions. Hence, to properly capture this consolidation of differing viewpoints, annotators are also guided to align review mentions that either sentimentally entail or contradict the summary aspect. For example, the two aforementioned conflicting review spans should be aligned to the summary span “the service was mostly good”.
Finally, annotators may mark multiple spans in reviews that redundantly represent the same statement. The guidelines also address paraphrasing, non-consecutive highlights, and unalignable summary spans. A detailed explanation of these guidelines can be found in Appendix A.2.
@@ 4.3 Annotator Training
The requirements of the aforementioned annotation process call for proficient-level annotations, which we achieved by means of controlled crowdsourcing Roit et al. (2020). We identified qualified annotators through three open qualification rounds, followed by three closed rounds for selected annotators, focusing on further training and refinement. Each open round involved annotators reading a brief task description and accordingly aligning information between a single summary sentence and a short review, on a simplified interface. After each open round, we reviewed the alignment and provided feedback. We then checked whether the annotators implemented our feedback in the following round (with a different sentence-review instance). If the annotators satisfyingly cooperated throughout the open rounds, they moved on to the closed rounds. Before the closed rounds, the qualified workers were asked to watch a 25-minute tutorial on the full annotation tool and guidelines (§4.1 and §4.2). The closed rounds were conducted similarly to the open rounds, but with a whole summary and review, with all guidelines, and on the full interface. The qualification process was fully compensated with a customary wage, requiring up to 5 minutes per round. From this process, we were able to gather 8 trained annotators, who annotated the 1000 instances in our dataset.
@@ 5.3 Trained Coverage Metric
Inspired by recent work that evaluates faithfulness and factuality using a dedicated trained model (Yin et al., 2021; Utama et al., 2022; Gekhman et al., 2023; Soleimani et al., 2023), we finetune an LLM that is tasked to assess whether the generated passage fully covers the highlights. In our methodology, each highlight is individually input along with the entire output, and the model outputs a binary answer for whether the highlight is contained in the passage. [^7] We also tried concatenating all the highlights together, and found it to be inferior. We derive synthesized training data for this task from our FiC dataset, using highlights and their corresponding summaries. For negative samples, we remove the summary sentence that aligns with the highlight. For positive samples, a random non-aligning summary sentence is omitted (to avoid a potential bias caused by sentence exclusion in the negative samples). We finetune a flan-t5-large model (Chung et al., 2022) with the synthesized coverage data. The input to the model is the highlight and modified summary, and the output is ‘yes’ or ‘no’, for positive and negative samples, respectively. The final score is the average probability of the token ‘yes’ across all highlights. [^8] We also explored an NLI-based coverage metric, where the passage serves as the premise and the highlights function as the hypothesis. We found it to achieve comparable results, however it requires substantially more computation time and memory. For more details, see Appendix E.
@@ 6.1 Experimental Setup
We examine several baseline models for solving the FiC task. The input to a model is a document set with spans marked within the documents (highlights), and the model is trained to generate a fused passage around the highlights.
@@ 6.2 Results
We apply our evaluation metrics on the proposed systems, with results presented in Table 3. We first observe that the exclusion of context from the input (Flan-T5only-H) yields the strongest faithfulness and coverage scores, yet the lowest coherence and redundancy scores. This shows the importance of incorporating context for more seamless outputs. Meanwhile, the removal of highlights (Flan-T5no-H) leads to a substantial degradation in faithfulness and coverage. This indicates that the Flan-T5H model indeed succeeds in learning to adjust the output according to the highlights, underlining the highlights’ role in enhancing the model’s performance.
Interestingly, even though the RL reward functions used in the RL-enriched model are the faithfulness and coverage metrics themselves, the outputs are eventually negatively affected when evaluating with these metrics. This result calls for a more in-depth investigation of enhanced reward functions that can leverage the benefits of RL training, as was shown to be helpful in Slobodkin et al. (2023a) for the single-input setup. We also find that single-shot GPT-4 yields the most coherent and least redundant texts. While it ranks highly in faithfulness and coverage, it is still overtaken by the finetuned Flan-T5only-H. Overall, our findings invite for further research on the FiC task, to develop fusion strategies that ensure comprehensive coverage and faithfulness to highlighted content, with coherent and low-redundancy outputs.
@@ 7 Conclusion
In this paper, we further promote the decomposition of grounded text generation as presented in (Slobodkin et al., 2022), extending it to the multi-document setting. To that end, we introduce the Fusion-in-Context (FiC) task, an extension of the task from (Slobodkin et al., 2022) which focuses on the content fusion step, to the multi-document setting. The FiC setting facilitates employing a single general-purpose fusion model for diverse content selection needs, capturing the challenges of repetitiveness and contradictions in source documents. It also supports interactive, user-driven generation, by allowing users to choose personalized content for the FiC module to merge into a customized passage. Moreover, direct access to the pre-selected ‘‘highlights’’ can facilitate attributed generation, where the pre-selected segments also serve as supporting cited content for the fused text. To advance the task, we introduce the FuseReviews[Uncaptioned image] benchmark, which includes a high-quality dataset, an evaluation framework for faithfulness and coverage of selected spans, and several baseline models to stimulate further research and exploration.
@@ Appendix B NLI Zero-Shot Prompt
Table 4: Faithfulness, coverage, and F-1 scores of the zero-shot GPT-4 model on 30 instances for the FiC development set, for varying numbers of in-context examples in the prompt. For each metric, the best scores are in bold.
@@ Appendix C GPT-4 Prompting
Table 4 presents the faithfulness, coverage, and F-1 scores of the zero-shot GPT-4 model across 30 instances from the FiC development set, for varying numbers of in-context examples in the prompt. Based on these outcomes, we chose to proceed with a single in-context example.
@@ Appendix D Fluency and Redundancy Human Annotation Protocol
We ask crowd-workers to assess the fluency and redundancy of the texts produced by all models under examination. We employ annotators who have demonstrated proficiency in semantic tasks, including summarization, in previous experiments. For evaluation purposes, 100 instances are randomly selected from our test set, and the texts generated by each model for these instances are evaluated, resulting in 500 total samples. Each sample is reviewed by three different annotators, and their scores are averaged to obtain a final assessment. The evaluation is facilitated through two Amazon Mechanical Turk interfaces, specifically designed for this study. One interface focuses on evaluating coherence, while the other assesses redundancy, with each interface presenting the annotators with one of the 500 samples (as depicted in Figure 4). Consistent with the methodology of (Slobodkin et al., 2022), a 5-point Likert scale is employed to rate the fluency and redundancy of the generated summaries. To minimize ambiguity and promote consistent ratings, each score on the scale is accompanied by explicit criteria (also illustrated in Figure 4). Taking into account an average response time of 30 seconds for each evaluation, we set the compensation for each response at 10 ¢.
@@
Figure 4: Example of the data collection interfaces used by the crowd-workers to evaluate the fluency (4(a)) and redundancy (4(b)) of summaries.
@@ E.3 Additional Meta Evaluation Setup Details
To achieve further agreement between the three authors, an additional reconciliation procedure was undertaken for the ten instances annotated by all three authors. This procedure entailed discussions for each instance where the annotations diverged by more than one point, separately for the faithfulness and coverage scores. During these discussions, each author explained the rationale behind their assigned score. Subsequently, the authors endeavored to reach a unanimous agreement on each instance, thereby further aligning their scoring criteria.